# Out-of-Distribution Detection in Molecular Complexes via Diffusion Models for Irregular Graphs

### v1

@@ 2.2 Bioinformatic Validation of OOD Log-Likelihoods
Figure 2: Log-likelihood distributions align with protein and ligand similarity metrics: Comparison of the distributions of a) log-likelihoods assigned by the diffusion model with the distributions of train-test similarity scores (b, c and d). b) Ligand Similarity: Calculated using Tanimoto scores between count-based molecular fingerprints. High scores indicate high similarity (1.0 = identical). c) Protein Similarity: Determined by TM-align based on optimal 3D protein structure alignment. High scores indicate high structural similarity (1.0 = identical) d) Aggregated Similarity: A composite score S=max(Tanimoto+TMScore+(1RMSD),0)S=max(Tanimoto+TMScore+(1-RMSD),0) was calculated as the sum of Tanimoto similarity, TM-scores, and inverted pocket-aligned ligand root mean squared deviation (RMSD). High scores signify highly similar protein-ligand complexes (3.0 = identical complexes). Each box (b, c and d) represents the distribution of $N=10,510$ similarity scores. Specifically, these are the scores between each training dataset complex and its most similar counterpart in the respective test dataset, as measured by that specific metric. Boxplots show the median (centre line), 25th–75th percentiles (box), whiskers extend to data points within 1.5 × IQR, outliers are not shown.
Across the eight datasets evaluated, the log-likelihood values assigned by the diffusion model align closely with the bioinformatic similarity estimates. This validation suggests that the PF-ODE log-likelihood distributions have a basis in the underlying structural and chemical similarity between the training and non-training datasets and that log-likelihood values are indicative of the degree of out-of-distribution of new data. However, despite this alignment, the log-likelihood distributions retained significant overlap with the training data, making robust separation by a classifier difficult. Consequently, we concluded that log-likelihoods alone are insufficient for reliable OOD detection. To address this, we incorporated additional PF-ODE trajectory statistics designed to capture the diffusion model’s struggle to process OOD data.
2.3 OOD Classification Using Trajectory Features
@@
We introduce a novel methodology for OOD detection that leverages the full information in a sample’s noising trajectories for accurate OOD classification. Instead of relying solely on the final scalar log-likelihood, we collect 18 additional trajectory statistics that characterize the model’s dynamical behavior when transforming an input structure into Gaussian noise.
@@
These additional features show coordinated variation under distribution shift caused by OOD samples: Stiff, bursty vector fields, higher acceleration and longer, less efficient paths tend to co-occur when processing OOD samples. Therefore, including quantities describing this coordinated movement facilitates robust OOD classification. We showcase this trajectory-aware OOD-detection approach using our protein-ligand datasets.
Classifier Setup
For each molecular complex, we collect a scalar log-likelihood and the 18 additional trajectory statistics, forming a $d=19$-dimensional feature vector describing the complexity, stability, and efficiency of the flow dynamics. To obtain an OOD detector, this set of trajectory features is subsequently used to fit separate Gaussian Kernel Density Estimators (KDEs) for in-distribution (ID) and OOD samples. To obtain an OOD score for new data, we derive a negated Log-Density Ratio (LDR), where larger scores indicate higher confidence that the sample is OOD. A binary classification is achieved by thresholding, optimized on a validation set (see Methods). To obtain a robust and balanced performance estimate despite class imbalance, we employ a bootstrap evaluation strategy. For each OOD split we partition ID and OOD complexes into training, validation, and test subsets. The PCA and the two KDEs are fit on their respective training data, and the decision threshold is fixed once by maximizing the F1-score for the OOD class on the validation set. Because the number of ID decoys can be much larger than the number of OOD binders, naive evaluation on the full test pool would overweight specificity. To obtain a balanced and robust estimate, we keep the OOD test set fixed and repeatedly draw balanced ID test subsets of the same size. For each draw we evaluate the classifier, treating OOD as the positive class, and record AUROC, accuracy, precision, recall, F1 and specificity.
Performance Evaluation
The classifier demonstrated robust performance on samples with strong distribution shifts relative to the training data (Table 1 and S3), achieving consistent separation on the most structurally unique OOD data such as HIV-proteases (3o9i) and heat-shock-proteins (2vw5). However, as anticipated, accuracy diminished on datasets with intermediate OOD characteristics, confirming that these datasets contain complexes that retained moderate similarity to the training data and are therefore hard to separate. A prime example is the CASF2016 benchmark: Originally sampled from the training distribution, this dataset was rendered more independent via CleanSplit filtering. Consequently, it occupies a OOD space between the training data and the OOD datasets, as confirmed by both log-likelihood profiles and bioinformatic similarity analysis (Figure 2). This moderate distribution shift resulted in the low OOD-classification accuracy. Similarity, the transporter protein dataset (3f3e) was predicted with low accuracy. This dataset equally showed intermediate OOD characteristics. Its log-likelihood distribution was very broad with significant overlap with the training log-likelihoods, a finding that was further confirmed with bioinformatic metrics showing a significant degree of retained structural similarity with the training data (Figure 2).
To assess the contribution of the trajectory features to the strong discriminative power of the OOD detector, we computed feature importance scores (Figure S2). While the PF-ODE log-likelihoods was most important for classification, we found a significant contribution of the additional trajectory descriptors to the OOD detector’s performance – the precise details are fleshed out in Section S1.3. Especially path stability, vector field burstiness and drift smoothness notably contribute to the OOD classification performance. Compared to a classifier based on PF-ODE log-likelihoods only, the classifier integrating additional trajectory descriptors achieved significantly more accurate OOD classification.
Trajectory features rescue 3dd0 classification
@@
One of the nine non-training datasets (3dd0, containing alpha-carbonic anhydrases) yielded unexpectedly high PF-ODE log-likelihoods. Relying on these scores alone would result in the erroneous classification of this dataset as in-distribution (Figure S4). However, our results demonstrate that augmenting the log-likelihoods with additional trajectory statistics sufficiently enriches the feature space to correct this prediction and flag the dataset as OOD. This ensemble of trajectory features enabled the construction of a robust decision boundary between the training data and the alpha-carbonic anhydrase structures, effectively compensating for the lack of discriminative signal in the raw likelihood scores.
@@
These observations made on our protein-ligand data show that modeling the joint behavior of several trajectory features yields substantially stronger OOD detection accuracy compared to methods that rely solely on the trajectory log-likelihoods. This confirms that these trajectory statistics provide critical, complementary signals that are characteristic fingerprints of out-of-distribution processing. While we demonstrated the power of this approach using protein–ligand complexes, a challenging domain of unordered 3D graphs, the implications extend far beyond molecular biology. This detection framework, using a set of 19 trajectory features, is universally applicable to any data modality amenable to continuous diffusion modeling. On top, due to the purely generative training objective, no data labels are required.
2.4 Correlation with GEMS errors
To assess whether log-likelihoods can predict the performance of predictive models, the binding affinity prediction model GEMS was trained on the same data split. This Graph Neural Network (GNN) processes graph representations of protein-ligand interaction patterns within a complex and performs regression to predict a scalar binding affinity value ($pK$). Although the diffusion model and GEMS are independent, the log-likelihoods assigned by the diffusion model to the training, validation, and test datasets generally align with the performance metrics achieved by GEMS on the same datasets. The binding affinity prediction $R^{2}$ of GEMS showed a strong correlation with the median log-likelihoods of the same dataset (Pearson Correlation Coefficient, $PCC=0.750$). Similar correlations were observed for the ranking performance of GEMS, with Kendall (Tau, $PCC=0.612$) and Spearman Correlation (Rho, $PCC=0.613$) coefficients correlating significantly with log-likelihoods (Figure 3).
Across the eight datasets evaluated, the mean binding affinity prediction errors of GEMS significantly correlated with the median log-likelihoods assigned by the diffusion model to those same datasets (PCC=0.880PCC=-0.880). Datasets assigned high median log-likelihoods consequently exhibit low mean prediction errors. Conversely, lower log-likelihood distributions are associated with a very steep increase in GEMS prediction errors (Figure 4). Regarding complex-wise errors, complexes with low log-likelihoods tend to be predicted with large errors by GEMS, while those with higher likelihoods yield lower errors. To describe this relationship, we performed a hierarchical fitting of five exponential curves (main fit, above fit, above-above fit, below fit, and below-below fit) to the log-likelihood-error relationship across a fixed number of samples drawn from all non-training distributions. This fitting confirmed that the relationship follows an approximate exponential behavior (Figure S1), showing that complex-wise log-likelihoods provide a rough estimate of the expected error for GEMS binding affinity predictions.
Figure 3: Correlation between log-likelihoods and GEMS performance: Comparison of log-likelihood distributions obtained for training, validation, CASF2016 and out-of-distribution (OOD) datasets with performance metrics achieved by the GEMS binding affinity prediction model trained on the same data. The box plot shows the distribution of log-likelihoods assigned to each dataset’s protein-ligand complexes by the diffusion model, where lower values signify increased deviation from the learned distribution (OOD) and higher values indicate high in-distribution probability (IID). The bar plot shows the corresponding performance metrics that GEMS achieved on the same datasets, including R-squared (R2), Kendall (Tau) and Spearman (Rho) rank correlation coefficients. Boxplots show the median (centre line), 25th–75th percentiles (box), whiskers extend to data points within 1.5 × IQR, outliers are not shown.
Figure 4: Out-of-distribution complexes yield higher GEMS errors: Heatmaps illustrating the relationship between the log-likelihoods assigned by the diffusion model and the corresponding GEMS binding affinity prediction errors (y-axis) across the training, validation, CASF2016 and out-of-distribution (OOD) datasets. The x-axis represents the log-likelihood for each complex, where lower values signify increased deviation from the learned distribution (OOD) and higher values indicate high in-distribution probability (ID). The y-axis represents the variance-normalized errors of the GEMS binding affinity prediction. Color intensity represents density, with white areas containing no data. To allow fair comparison across datasets with different label spreads, absolute errors are normalized by the label variance, preventing artificially low error values on narrowly distributed datasets. The heatmaps visually confirm that low log-likelihoods are associated with higher GEMS prediction errors.
PF-ODE likelihood as a typicality and risk signal
We interpret the PF-ODE negative log-likelihood $L(x)$ as a task-agnostic typicality score under the diffusion density $q_{\phi}$ learned from in-distribution (ID) complexes.
$L(x)\;:=\;-\log q_{\phi}(x),$
Empirically, this typicality is also predictive of downstream reliability: complexes with unusually large $L(x)$ (low likelihood under $q_{\phi}$) tend to incur larger binding-affinity errors under GEMS (Figs. 4S1). The key intuition is that if prediction error grows (roughly) monotonically as inputs become less typical, then concentration of $L(x)$ under the ID law immediately yields a high-probability error guarantee. This is the content of the following informal proposition:
Proposition 2.1 (Likelihood controls error with high probability).
2.5 GEMS-based OOD Detection Baselines
@@
To rigorously benchmark our trajectory-aware OOD detection method, we compared it against a family of classical anomaly detectors operating on the fixed latent representations of the pretrained GEMS encoder. We treated the GEMS encoder as a static feature extractor and assumed that ID complexes occupy a relatively compact region of the GEMS representation space, and that OOD complexes are either geometrically distant from this region or sparsely supported. For every molecular complex, we computed a summary embedding by averaging the output vectors from multiple stochastic forward passes of GEMS. We then trained five unsupervised anomaly detectors on these standardized ID embeddings, covering distance–based, density–ratio, isolation–based and margin–based approaches:
@@
To benchmark our method against state-of-the-art uncertainty estimation techniques, we implemented Rate-In [10], a dynamic baseline that is built on the concept of Monte Carlo dropout. Instead of a fixed global dropout rate, Rate-In applies adaptive, per-sample, per-layer dropout rates chosen to preserve information flow through the network. It works under the assumption that ID samples allow for relatively high dropout rates without disrupting the computation, whereas OOD inputs are more fragile and loose information under dropout perturbations.
We integrated this adaptive scheme directly into the GEMS binding affinity predictor. For each molecular complex, the method iteratively optimizes the dropout intensity at every network layer, seeking the maximum dropout level that still preserves stable information flow. This process generates a sample-specific robustness profile. The final anomaly score is constructed by aggregating three distinct signals: The standard predictive variance (epistemic uncertainty), the heterogeneity of the optimized dropout rates across layers, and the computational effort (convergence difficulty) required to find these stable rates. High instability in these metrics indicates that the predictive model struggles to process the input structure, suggesting the input is OOD.
@@
These baselines establish a rigorous reference for evaluating our trajectory-aware method. While Rate-In provides a sophisticated measure of predictive uncertainty, our empirical analysis showed that it offered lower discrimination accuracy compared to the trajectory–aware OOD classifier approach (Table S4). This demonstrates that probing the stability of a discriminative model is inferior to the sensitivity of the diffusion-based approach using multiple PF-ODE trajectory features. Similarly, the embedding space baseline showed lower OOD detection accuracy across most datasets, showing that the latent space features from the GEMS encoder possess less OOD separability than the descriptors extracted from the PF-ODE trajectory.
@@
Table 2: Comparison of AUROC (mean $\pm$ standard deviation) across OOD datasets for the trajectory-aware LDR diffusion-based detector (Traj-LDR), the embedding-space GEMS baseline, and the Rate-In baseline. All values are taken from Tables S3, S2, and S4. The validation split used only for Rate-In is omitted here.
@@
The bioinformatic similarity analysis (Sections 2.2) and the GEMS-based OOD detection baselines (Section 5.6) provide a strong reference for our trajectory-aware OOD detection. However, our diffusion-based OOD detector occupies a fundamentally different region in the design space than the deterministic baselines. The diffusion model is trained in a purely generative fashion. It only observes bare three-dimensional coordinates and discrete atom/residue identities, and its objective is to approximate the joint data distribution. It never sees binding affinity labels, similarity scores, or any task-specific supervision. By contrast, out baselines are heavily endowed with biochemical priors: The bioinformatic OOD approximations are hand-crafted around well-established notions of structural and chemical similarity like TM-scores, Tanimoto scores and ligand RMSD values. In the Embedding Space baseline, OOD decisions are based on a static snapshot of the GEMS representation geometry. Similarly, the Rate-In baseline makes use of the pretrained weights of the GEMS model and derives OOD scores by probing prediction stability under dropout perturbation.
In this sense, the diffusion model is substantially more structure-agnostic than the baselines, yet it is asked to solve a more demanding problem (full generative modelling) from strictly weaker supervision. Nevertheless, the OOD detection with diffusion trajectory features outperforms all baselines in our task of recognizing OOD datasets of isolated protein families (Table 2).
@@
3 Conclusions
We introduced a diffusion-based framework for quantifying out-of-distribution (OOD) structure in irregular 3D graph data, instantiated on protein–ligand complexes. Reliable deployment of predictive models in structural biology and drug discovery requires knowing when a new complex lies outside the effective support of the training data. Our central finding is that a generative model trained only on raw 3D geometry and discrete biochemical types can provide a principled, task-agnostic OOD signal in its dynamic noising trajectories, and that this signal is strongly informative about the reliability of an independent binding-affinity predictor.
Methodologically, we use a unified continuous diffusion process that treats 3D coordinates and discrete atom/residue identities in a single Euclidean state, with categorical types embedded on an L2-normalised sphere. Posterior–mean interpolation turns logits into continuous scores, yielding a stable SE(3)-equivariant denoiser whose probability–flow ODE (PF–ODE) defines per-complex log-likelihoods. These PF–ODE likelihoods align well with bioinformatic similarity measures. Moreover, they correlate tightly with GEMS performance across datasets and follow a calibrated envelope of complex-wise prediction errors, enabling a priori risk estimates for downstream models.
Beyond scalar likelihoods, we show that trajectory-level statistics of the PF–ODE, capturing path geometry, stiffness, vector-field spikiness, energetic cost, and drift smoothness—carry complementary OOD information. A simple density-ratio classifier in this feature space outperforms strong non-diffusion baselines built on GEMS embeddings and adaptive uncertainty (Rate-In), despite never using binding-affinity labels.
This trajectory-aware OOD detection approach can offer a computational advantage: Given a diffusion model trained on the training distribution, trajectory features are fast to obtain and can potentially replace slower and more computationally intensive deterministic methods of OOD quantification (such as those described above), especially in settings where new data appears frequently and OOD assessment must be performed rapidly. Moreover, for other irregular graph data, robust complementary methods for deterministic OOD quantification may not exist. In these cases, likelihoods assigned by a diffusion model might be the only viable method for estimating a sample’s OOD degree.
Because the framework is fully generative and label-free, it extends beyond protein–ligand complexes to any modality that admits continuous diffusion, offering a general recipe for OOD quantification and error forecasting in geometric deep learning.

### v2

@@ 2.3 Log-Likelihoods Align with Bioinformatic OOD Levels
Figure 4: Log-likelihood distributions align with bioinformatic similarity metrics: a) Distributions of log-likelihoods assigned by the diffusion model to the validation set (N=949) and to the test datasets 1nvq (N=2560), 1sqa (N=671), 2p15 (N=412), 2vw5 (N=204), 3dd0 (N=454), 3f3e (N=233), 3o9i (N=463) and CASF2016 (N=272). b) The distributions of similarity scores S=max(Tanimoto+TMScore+(1RMSD),0)S=max(Tanimoto+TMScore+(1-RMSD),0) between the training dataset and the test datasets. Scores are calculated as the complex-wise sum of Tanimoto similarity, TM-scores, and inverted pocket-aligned ligand root mean squared deviation (RMSD). High scores signify highly similar protein-ligand complexes (3.0 = identical). Each box represents the distribution of $N=10,510$ similarity scores. Specifically, these are the scores between each training dataset complex and its most similar counterpart in the respective test dataset, as measured by that specific metric. Boxplots show the median (centre line), 25th–75th percentiles (box), whiskers extend to data points within 1.5 × IQR, outliers are not shown.
2.4 Log-Likelihoods are Predictive of GEMS Errors
A critical application of OOD detection is to preemptively flag samples where predictive models are likely to fail. If the diffusion model’s PF-ODE log-likelihoods accurately reflect the degree of data unfamiliarity, they should predict the performance degradation of independent models trained on the same dataset. This would allow the diffusion model to act as a quality filter, warning users when a new sample is too novel for reliable prediction. To test this capability, we trained the GEMS binding affinity model on the same dataset. This graph neural network processes graph representations of protein-ligand interactions and predicts a scalar binding affinity value. Using the test datasets, we analyzed whether low diffusion likelihoods consistently correspond to high prediction errors.
Across the eight evaluated datasets, we observed a strong correlation between the diffusion model’s median log-likelihoods and the predictive performance of GEMS on the same datasets. This relationship was significant both in terms of the coefficient of determination ($R^{2}$, Pearson Correlation Coefficient $r=0.750$) and the mean absolute error (r=0.880r=-0.880). High-likelihood (ID) datasets consistently yielded more accurate predictions (Figure 5), while low likelihood distributions (OOD) showed a steep increase in error (Figure 6). At the individual complex level, this trend follows an approximate exponential behavior: Samples with low likelihoods are prone to large prediction errors, whereas complexes with high likelihoods are predicted accurately with high probability (Figure S1).
Figure 5: Correlation between log-likelihoods and GEMS performance: Comparison of log-likelihood distributions obtained for training, validation, CASF2016 and out-of-distribution (OOD) datasets with performance metrics achieved by the GEMS binding affinity prediction model trained on the same data. The box plot shows the distribution of log-likelihoods assigned to each dataset’s protein-ligand complexes by the diffusion model, where lower values signify increased deviation from the learned distribution (OOD) and higher values indicate high in-distribution (ID) probability. The bar plot shows the corresponding performance metrics that GEMS achieved on the same datasets, including R-squared (R2), Kendall (Tau) and Spearman (Rho) rank correlation coefficients. Boxplots show the median (centre line), 25th–75th percentiles (box), whiskers extend to data points within 1.5 × IQR, outliers are not shown.
Figure 6: Out-of-distribution complexes yield higher GEMS errors: Heatmaps illustrating the relationship between the log-likelihoods assigned by the diffusion model and the corresponding GEMS binding affinity prediction errors (y-axis) across the training, validation, CASF2016 and out-of-distribution (OOD) datasets. The x-axis represents the log-likelihood for each complex, where lower values signify increased deviation from the learned distribution (OOD) and higher values indicate high in-distribution (ID) probability. The y-axis represents the variance-normalized errors of the GEMS binding affinity prediction. Color intensity represents density, with white areas containing no data. To allow fair comparison across datasets with different label spreads, absolute errors are normalized by the label variance, preventing artificially low error values on narrowly distributed datasets. The heatmaps visually confirm that low log-likelihoods are associated with higher GEMS prediction errors.
We interpret the PF-ODE negative log-likelihood $L(x)$ as a task-agnostic typicality score under the diffusion density $q_{\phi}$ learned from ID complexes.
$L(x)\;:=\;-\log q_{\phi}(x),$
Empirically, this typicality is also predictive of downstream reliability: Complexes with unusually large $L(x)$ (low likelihood under $q_{\phi}$) tend to incur larger binding-affinity errors under GEMS (Figures 6 and S1). The key intuition is that if prediction error grows (roughly) monotonically as inputs become less typical, then concentration of $L(x)$ under the ID law immediately yields a high-probability error guarantee. This is the content of the following informal proposition:
Proposition 2.1 (Likelihood controls error with high probability).
2.5 Robust OOD Classification via Trajectory Features
@@
The strong correlation between diffusion log-likelihoods and bioinformatic similarity across eight test datasets suggests that the PF-ODE log-likelihoods are rooted in the underlying chemical similarity. Nevertheless, they prove insufficient for robust OOD detection in isolation. The log-likelihood distributions exhibit significant overlap with the training data, making accurate classification difficult. Moreover, the metric fails completely on low-complexity interactions due to complexity bias (3dd0, Figure 3h). To resolve these limitations and improve separability, we introduce a novel methodology for OOD detection that leverages the full information in a sample’s PF-ODE trajectories for accurate classification. Instead of relying solely on the final scalar log-likelihood, we collect 18 additional PF-ODE trajectory statistics that characterize the model’s dynamical behavior when transforming an input structure into Gaussian noise.
@@
These additional features show coordinated variation under distribution shift: Stiff, bursty vector fields, higher acceleration and longer, less efficient paths tend to co-occur when processing OOD samples. Therefore, including quantities describing this coordinated movement facilitates robust OOD classification. To build a classifier that synthesizes these 19-dimensional feature space into an OOD signal, we fit separate Gaussian Kernel Density Estimators (KDEs) to ID and OOD samples. For new data, we derive a negated Log-Density Ratio (LDR), where larger scores indicate higher OOD confidence. A binary classification is achieved by thresholding, optimized on a validation set (Section 4.5). We refer to this classifier as trajectory-aware LDR classifier.
Trajectory statistics overcome complexity bias
@@
One of the nine non-training datasets (3dd0, containing $\alpha$-carbonic anhydrases) yielded unexpectedly high PF-ODE log-likelihoods (Figure 3). Relying on these scores alone would result in the erroneous classification of this dataset as in-distribution. However, our results demonstrate that augmenting the log-likelihoods with additional trajectory statistics sufficiently enriches the feature space to correct this prediction and flag the dataset as OOD with high confidence. This ensemble of trajectory features enabled the construction of a robust decision boundary between the training data and the $\alpha$-carbonic anhydrase structures, effectively compensating for the lack of discriminative signal in the raw likelihood scores.
Classifier Performance
To ensure robust performance estimates, we employed a balanced bootstrap evaluation strategy for each test dataset. Because the number of ID decoys is much larger than the number of OOD binders, we keep the OOD test set fixed and repeatedly draw 100 balanced ID test subsets of the same size. For each draw we evaluate the classifier and record AUROC as a measure of ID/OOD separability. We observed robust performance on test datasets with strong distribution shifts relative to the training data (Table 1 and S3), achieving consistent separation on the most structurally unique OOD data such as HIV-proteases (3o9i), heat-shock-proteins (2vw5) and $\alpha$-carbonic anhydrases (3dd0). However, as anticipated, accuracy diminished on datasets with intermediate OOD characteristics, confirming that these datasets contain complexes that retained moderate similarity to the training data and are therefore hard to separate. A prime example is the CASF2016 benchmark: Originally sampled from the training distribution, this dataset was rendered more independent via CleanSplit filtering. Consequently, it occupies a OOD space between the training data and the OOD datasets, as confirmed by both log-likelihood profiles and bioinformatic similarity analysis (Figure 4). This moderate distribution shift resulted in the low OOD-classification accuracy. Similarity, the transporter protein dataset (3f3e) was predicted with low accuracy. This dataset equally showed intermediate OOD characteristics. Its log-likelihood distribution was very broad with significant overlap with the training log-likelihoods, a finding that was further confirmed with bioinformatic metrics showing a significant degree of retained structural similarity with the training data (Figure 4).
@@
Feature Importance
To understand which aspects of the PF-ODE trajectory drive separation between ID and OOD complexes, we quantify the importance of each of our $d=19$ trajectory statistics and rank them accordingly (Figure 7). Rather than reporting a single heuristic, we use a combined importance score that rewards features which separate ID and OOD strongly and do so consistently across complexes and bootstrap splits.
Several patterns stand out: First, PF-ODE log-likelihood is the single most discriminative coordinate, but it is not the whole story: Multiple trajectory descriptors achieve comparable importance, confirming that OOD shift manifests as a dynamic signature along the probability flow rather than only as a terminal density value. Second, the highest-ranked non-likelihood features cluster into interpretable families: (i) path geometry (tortuosity/efficiency), indicating that OOD inputs induce longer and less direct transport to noise; (ii) field magnitude and burstiness (vf_l2 mean/max/std and spikiness), indicating larger and more intermittent corrective updates; and (iii) stability and smoothness (Lipschitz and acceleration), indicating stiffer and less temporally regular dynamics. Finally, coupling-related terms (e.g. coupling_consistency and dynamic_coord_feature_coupling) rank non-trivially, suggesting that OOD behavior is not purely geometric: Atypical complexes also perturb how coordinate and identity updates co-evolve across noise scales.
Figure 7: Ranking of PF-ODE trajectory features by discriminative power for OOD detection: Relative importance of 19 trajectory statistics, quantified by a composite score of separation strength (effect size between ID/OOD distributions) and stability under resampling. While log-likelihood is the dominant feature, dynamic characteristics such as path tortuosity, vector field energy, and Lipschitz stability contribute significantly to detection performance. Color indicators classify features into interpretable families (e.g., Path Geometry, Field Magnitude, Stability, Feature-Coordinate Coupling and Structural Consistency). This ranking confirms that OOD samples exhibit distinct dynamic noising trajectories, traversing longer, more energetic, and less stable paths.
2.6 GEMS-based Post-Hoc Baselines
Due to the novelty of the 3D graph OOD detection task, there is currently a lack of competing models or established benchmarks to serve as baselines for our framework. Therefore, we constructed our own suite of post-hoc baselines by adapting two distinct paradigms from other machine learning domains. First, we implemented a family of classical anomaly detectors operating on the fixed latent representations of the pretrained GEMS encoder to test if OOD signals are captured in static feature spaces. Second, we implemented Rate-In, a state-of-the-art dynamic uncertainty baseline, which probes the stability of GEMS under dropout.
@@ 2.6 GEMS-based Post-Hoc Baselines
We treated the GEMS encoder as a static feature extractor and assumed that ID complexes occupy a relatively compact region of the GEMS representation space, and that OOD complexes are either geometrically distant from this region or sparsely supported. For every molecular complex, we computed a summary embedding by averaging the output vectors from multiple stochastic forward passes of GEMS. We then trained five unsupervised anomaly detectors on these standardized ID embeddings, covering distance–based, density–ratio, isolation–based and margin–based approaches:
@@
We implemented Rate-In [38], a dynamic baseline that is built on the concept of Monte Carlo dropout. Instead of a fixed global dropout rate, Rate-In applies adaptive, per-sample, per-layer dropout rates chosen to preserve information flow through the network. It works under the assumption that ID samples allow for relatively high dropout rates without disrupting the computation, whereas OOD inputs are more fragile and loose information under dropout perturbations. We integrated this adaptive scheme directly into the GEMS binding affinity predictor. For each molecular complex, the method iteratively optimizes the dropout intensity at every network layer, seeking the maximum dropout level that still preserves stable information flow. This process generates a sample-specific robustness profile. The final anomaly score is constructed by aggregating three distinct signals: The standard predictive variance (epistemic uncertainty), the heterogeneity of the optimized dropout rates across layers, and the computational effort (convergence difficulty) required to find these stable rates. High instability in these metrics indicates that the predictive model struggles to process the input structure, suggesting the input is OOD.
@@
These baselines establish a rigorous reference for evaluating our trajectory-aware method. While Rate-In provides a sophisticated measure of predictive uncertainty, our empirical analysis showed that it offered lower discrimination accuracy compared to the trajectory–aware OOD classifier approach (Table 2). This demonstrates that probing the stability of a discriminative model is inferior to the sensitivity of the diffusion-based approach using multiple PF-ODE trajectory features. Similarly, the embedding space baseline showed lower OOD detection accuracy across most datasets, showing that the latent space features from the GEMS encoder possess less OOD separability than the descriptors extracted from the PF-ODE trajectory. Interestingly, both the embedding-space baseline and Rate-In outperformed our method on the transporter protein dataset (3f3e). This suggests that for certain classes of distribution shift, the static latent representation of GEMS may capture specific OOD characteristics that are more discriminative than the PF-ODE flow dynamics of our diffusion model.
@@
Table 2: Comparison of AUROC (mean $\pm$ standard deviation) across OOD datasets for the trajectory-aware LDR diffusion-based detector (Traj-LDR), the embedding-space GEMS baseline, and the Rate-In baseline. All values are taken from Tables S3, S1 and S2. The validation split used only for Rate-In is omitted here.
@@
The bioinformatic similarity analysis (Sections 2.1) and the GEMS-based post-hoc OOD detection baselines (Section 2.6) provide a strong reference for our trajectory-aware OOD detection. However, our diffusion-based OOD detector occupies a fundamentally different region in the design space than the deterministic baselines. The diffusion model is trained in a purely generative fashion. It only observes bare three-dimensional coordinates and discrete atom/residue identities, and its objective is to approximate the joint data distribution. It never sees binding affinity labels, similarity scores, or any task-specific supervision. By contrast, out baselines are heavily endowed with biochemical priors: The bioinformatic OOD approximations are hand-crafted around well-established notions of structural and chemical similarity like TM-scores, Tanimoto scores and ligand RMSD values. In the Embedding Space baseline, OOD decisions are based on a static snapshot of the GEMS representation geometry. Similarly, the Rate-In baseline makes use of the pretrained weights of the GEMS model and derives OOD scores by probing prediction stability under dropout perturbation.
In this sense, the diffusion model is substantially more structure-agnostic than the baselines, yet it is asked to solve a more demanding problem (full generative modeling) from strictly weaker supervision. Nevertheless, the OOD detection with diffusion trajectory features outperforms all baselines in our task of recognizing OOD datasets of isolated protein families (Table 2).